A feather is dropped on the moon from a height of 1.40 meters. The acceleration of gravity on the moon is 1.67 m/s2. Determine the time for the feather to fall to the surface of the moon.
1
Expert's answer
2020-05-21T13:33:05-0400

The law of motion of the object falling from height "H"is "y(t) = H - \frac{g t^2}{2}", where "g" is the acceleration of the gravity. When the object hits the surface "y(t) = 0", from where "t = \sqrt{\frac{2 H}{g}}".

Substituting "H = 1.4 m" and "g = 1.67 \frac{m}{s^2}" into the last formula, obtain "t \approx 1.3 s".